# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

### v1

@@ D.1.2 Lemmata for the CCOD estimator
by Condition 2.2. Then, by notation in Condition 2.4, $I_{1}+I_{2}\leq\eta^{2}\{\bar{r}_{n,t,a,1}+\bar{r}_{n,t,a,2}+\bar{r}_{n,t,a,3}\}$. Since $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}\leq 2$,
@@
by Condition 2.4. This established the pointwise RAL property: $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{t,a})+o_{p}(n^{-1/2})$. Since $\varphi^{*\text{CCOD}}_{t,a}$ is uniformly bounded, $\mathbb{P}\{(\varphi^{*\text{CCOD}}_{t,a})^{2}\}<\infty$ and since $\mathbb{P}\{\varphi^{*\text{CCOD}}_{t,a}\}=0$, it follows that
@@
For the uniform RAL, we use the same decomposition above. By conditions 2.1, 2.2 and 2.4 and Lemma D.4,
@@
which is $o_{p}(n^{-1/2})$ by Condition 2.5. Therefore, $\sup_{u\in[0,t]}\left|\widehat{\theta}^{\text{CCOD}}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a})\right|=o_{p}(n^{-1/2})$. Since $\{\varphi^{*\text{CCOD}}_{u,a}:u\in[0,t]\}$ is a uniformly bounded $\mathbb{P}$-Donsker class by Lemma D.3, $\{n^{1/2}\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a}):u\in[0,t]\}$ converges weakly to a tight mean-zero Gaussian process with covariance $(u,v)\mapsto\mathbb{P}(\varphi^{*\text{CCOD}}_{u,a}\varphi^{*\text{CCOD}}_{v,a})$. ∎
@@
If we only focus on the consistency of $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$, then the condition $\bar{\pi}_{\infty}=\bar{\pi}$, $\bar{G}_{\infty}=\bar{G}$, $q^{0}_{\infty}=q^{0}$ and $\bar{S}_{\infty}=\bar{S}$ can be replaced by the following statement: For $\mathbb{P}$-almost all $\mathbf{x}$, there exist measurable sets $\bar{\mathcal{S}}_{x},\bar{\mathcal{G}}_{x}\subseteq[0,t]$ such that $\bar{\mathcal{S}}_{x}\cup\bar{\mathcal{G}}_{x}=[0,t]$ and $\bar{\Lambda}(u\mid a,\mathbf{x})=\bar{\Lambda}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\bar{\mathcal{S}}_{x}$ and $\bar{G}(u\mid a,\mathbf{x})=\bar{G}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\bar{\mathcal{G}}_{x}$. In addition, if $\bar{\mathcal{S}}_{x}$ is a strict subset of $[0,t]$, then $\bar{\pi}(a\mid\mathbf{x})=\bar{\pi}_{\infty}(a\mid\mathbf{x})$ and $q_{0}(\mathbf{x})=q^{0}_{\infty}(\mathbf{x})$ as well. Then, $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$ is consistent if Conditions 2.1 and 2.2 hold, and it is uniform consistent if Condition 2.3 also holds.
@@
The proof details of Remark D.6 are similar to those in the proof of Theorem 2 in Westling et al. (2023), which are omitted here. A sketch is that by decomposing the integral $\int_{0}^{t}$ as $\int_{\bar{\mathcal{S}}_{x}}+\int_{\bar{\mathcal{S}}^{c}_{x}}$, where $\bar{\mathcal{S}}^{c}_{x}$ is the complement of set $\bar{\mathcal{S}}_{x}$, and $\bar{\mathcal{S}}_{x}^{c}\subseteq\bar{\mathcal{G}}_{x}$ by definition. Then it is straightforward to verify that when the statement in Remark D.6 holds, the following integral (also in the bias term (4))
@@ D.2 Theory of the source-site estimator
Following a similar argument in the proof of Theorem 2.8 in Section D.1.1, we can express the integrand of the second term above as
@@
$\quad+\iiiint\mathbb{I}(u\leq t,u\leq y)\frac{S^{k}(t\mid a,\mathbf{x})S^{k}(u-\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid\mathbf{x})}\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)N^{k}_{1}(du\mid a,\mathbf{x})\mu(d\mathbf{x})$
@@
$\quad+\iiint S^{k}(t\mid a,\mathbf{x})\int_{0}^{t\wedge y}\frac{S^{k}(u-\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid\mathbf{x})^{2}}N^{k}_{1}(du\mid a,\mathbf{x})\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)\mu(d\mathbf{x})$
$=\mathbb{E}\left[S^{k}_{A}(t\mid\mathbf{X})\frac{I(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}\right\}\dot{\ell}(Y,\Delta\mid A,\mathbf{X},R=k)\right],$
@@
$\mathbb{E}\left[H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A=a,\mathbf{X}=\mathbf{x},R=k\right]=0$
@@
$=\mathbb{E}\left[S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k\mid\mathbf{X})}\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\right].$
@@
$\frac{\partial}{\partial\epsilon}\theta^{0}(t,a)\bigg{|}_{\epsilon=0}$ $=\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\mathbb{E}\bigg{\{}S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)\mathbb{I}(R=k)}{\pi^{k}(a\mid\mathbf{X})\mathbb{P}(R=k\mid\mathbf{X})}$
×$\quad\quad\times\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(y\mid\mathbf{X})G^{k}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})~{}\bigg{|}~{}\mathbf{X}\bigg{\}}\bigg{]}$
$=\mathbb{E}\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\{S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)\}\dot{\ell}(\mathcal{O})\right]-\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=0)}\frac{\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=k\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}$
×$\quad\quad\times\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(y\mid\mathbf{X})G^{k}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\bigg{]}.$
@@
$\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})$ $=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=0)\mathbb{P}(R=k\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
@@ D.2.2 Regularity conditions and RAL of the source-site estimator
If Conditions D.1D.2 hold, with $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$, the CCOD holds, and Condition D.4 also holds, then $\widehat{\theta}^{k,0}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*k,0}_{t,a})+o_{p}(n^{-1/2})$. In particular, $n^{1/2}(\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a))$ then convergences in distribution to a normal random variable with mean zero and variance $\sigma^{2}=\mathbb{P}[(\varphi^{*k,0}_{t,a})^{2}]$. If Conditions D.3 and D.5 also hold, then
@@ D.2.3 Lemmata for the source-site estimator
$\widehat{\theta}_{n}^{k}(t,a)-\theta^{0}(t,a)=\mathbb{P}_{n}[\varphi^{*k,0}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a}\right]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right].$ (6)
@@
Under Condition D.2, there exists a universal constant $C=C(\eta)$ such that for each $k$, $m$, $n$, $t$, and $a$,
@@
$B^{k}(\mathcal{V}_{m})$ $=\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right],$
@@
The expression of $\widehat{B}^{k}_{m}(\mathcal{V}_{m})-B^{k}_{\infty}(\mathcal{V}_{m})$ is exactly the same as the Lemma 3 in Westling et al. (2023), while we only need to replace the corresponding nuisance functions by the site-$k$ version here, so the detail is omitted. By the triangle inequality, we have $\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}-\varphi^{k,0}_{\infty,t,a}]^{2}\leq\left\{\sum_{j=1}^{4}\{\mathbb{P}[(U^{k}_{j,n,m,t,a})^{2}]\}^{1/2}\right\}^{2}$. Therefore, under Assumption 2.3 and Condition D.2, there exists a universal constant $C=C(\eta)$ such that the result in the statement holds. Thus, the proof is completed. ∎
@@
If Conditions D.1D.2 hold, then $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}^{1/2}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a_{0}}\right]=o_{p}(n^{-1/2})$. If Conditions D.3 holds as well,
@@
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}S^{k}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})}\left\{\frac{\omega^{k,0}_{\infty}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})\pi^{k}_{\infty}(a\mid\mathbf{X})}-1\right\}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\right].$
@@
$\mathbb{E}\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\{S^{0}_{\infty}(t\mid a,\mathbf{X})-S^{0}(t\mid a,\mathbf{X})\}\right]$
$\quad+\mathbb{E}\left[\frac{q^{k}(\mathbf{X})}{\mathbb{P}(R=k)}\omega^{k,0}_{\infty}(\mathbf{X})S^{k}_{\infty}(t\mid a,\mathbf{X})\frac{\pi^{k}(a\mid\mathbf{X})}{\pi^{k}_{\infty}(a\mid\mathbf{X})}\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})G^{k}(y\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\right]$
$=\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\{S^{0}_{\infty}(t\mid a,\mathbf{X})-S^{0}(t\mid a,\mathbf{X})\}\right]$
$\quad+\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\omega^{k,0}_{\infty}(\mathbf{X})}{\omega^{k,0}(\mathbf{X})}S^{k}_{\infty}(t\mid a,\mathbf{X})\frac{\pi^{k}(a\mid\mathbf{X})}{\pi^{k}_{\infty}(a\mid\mathbf{X})}\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})G^{k}(y\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\right].$

### v2

@@ D.1.3 Lemmata for the CCOD estimator
by Condition D.2. Then, by notation in Condition D.3, $I_{1}+I_{2}\leq\eta^{2}\{\bar{r}_{n,t,a,1}+\bar{r}_{n,t,a,2}+\bar{r}_{n,t,a,3}\}$. Since $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}\leq 2$,
@@
by Condition D.3. This established the pointwise RAL property: $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{t,a})+o_{p}(n^{-1/2})$. Since $\varphi^{*\text{CCOD}}_{t,a}$ is uniformly bounded, $\mathbb{P}\{(\varphi^{*\text{CCOD}}_{t,a})^{2}\}<\infty$ and since $\mathbb{P}\{\varphi^{*\text{CCOD}}_{t,a}\}=0$, it follows that
@@
For the uniform RAL, we use the same decomposition above. By conditions D.1, D.2 and D.3 and Lemma D.4,
@@
which is $o_{p}(n^{-1/2})$ by Condition D.5. Thus, $\sup_{u\in[0,t]}\left|\widehat{\theta}^{\text{CCOD}}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a})\right|=o_{p}(n^{-1/2})$. Since $\{\varphi^{*\text{CCOD}}_{u,a}:u\in[0,t]\}$ is a uniformly bounded $\mathbb{P}$-Donsker class by Lemma D.3, $\{n^{1/2}\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a}):u\in[0,t]\}$ converges weakly to a tight mean-zero Gaussian process with covariance $(u,v)\mapsto\mathbb{P}(\varphi^{*\text{CCOD}}_{u,a}\varphi^{*\text{CCOD}}_{v,a})$. ∎
@@
If we only focus on the consistency of $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$, then the condition $\bar{\pi}_{\infty}=\bar{\pi}$, $\bar{G}_{\infty}=\bar{G}$, $q^{0}_{\infty}=q^{0}$ and $\bar{S}_{\infty}=\bar{S}$ can be replaced by the following statement: For $\mathbb{P}$-almost all $\mathbf{x}$, there exist measurable sets $\bar{\mathcal{S}}_{x},\bar{\mathcal{G}}_{x}\subseteq[0,t]$ such that $\bar{\mathcal{S}}_{x}\cup\bar{\mathcal{G}}_{x}=[0,t]$ and $\bar{\Lambda}(u\mid a,\mathbf{x})=\bar{\Lambda}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\bar{\mathcal{S}}_{x}$ and $\bar{G}(u\mid a,\mathbf{x})=\bar{G}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\bar{\mathcal{G}}_{x}$. In addition, if $\bar{\mathcal{S}}_{x}$ is a strict subset of $[0,t]$, then $\bar{\pi}(a\mid\mathbf{x})=\bar{\pi}_{\infty}(a\mid\mathbf{x})$ and $q_{0}(\mathbf{x})=q^{0}_{\infty}(\mathbf{x})$ as well. Then, $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$ is consistent if Conditions D.1 and D.2 hold, and it is uniform consistent if Condition D.4 also holds.
@@
The proof details of Remark D.6 are similar to those in the proof of Theorem 2 in Westling et al. [48], which are omitted here. A sketch is that by decomposing the integral $\int_{0}^{t}$ as $\int_{\bar{\mathcal{S}}_{x}}+\int_{\bar{\mathcal{S}}^{c}_{x}}$, where $\bar{\mathcal{S}}^{c}_{x}$ is the complement of set $\bar{\mathcal{S}}_{x}$, and $\bar{\mathcal{S}}_{x}^{c}\subseteq\bar{\mathcal{G}}_{x}$ by definition. Then it is straightforward to verify that when the statement in Remark D.6 holds, the following integral (also in the bias term (D.1.3))
@@ D.2 Theory of the source-site estimator
Following a similar argument in the proof of Theorem 2.7 in Section D.1.1 (i.e., the way we derive the derivative for $\bar{S}_{\epsilon}$ wrt $\epsilon$ there), we can express the integrand of the second term above as
@@
$\quad+\iiiint\mathbb{I}(u\leq t,u\leq y)\frac{S^{k}(t\mid a,\mathbf{x})S^{k}(u-\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid\mathbf{x})}$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)N^{k}_{1}(du\mid a,\mathbf{x})\mu(d\mathbf{x})$
@@
$\quad+\iiint S^{k}(t\mid a,\mathbf{x})\int_{0}^{t\wedge y}\frac{S^{k}(u-\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid\mathbf{x})^{2}}N^{k}_{1}(du\mid a,\mathbf{x})$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)\mu(d\mathbf{x})$
$=\mathbb{E}\bigg{[}S^{k}(t\mid a,\mathbf{X})\frac{I(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{H^{k}(t\wedge Y,a,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid a,\mathbf{X})}{S^{k}(Y\mid a,\mathbf{X})D^{k}(Y\mid a,\mathbf{X})}\right\}$
×˙$\qquad\qquad\times\dot{\ell}(Y,\Delta\mid a,\mathbf{X},R=k)\bigg{]},$
@@
$\mathbb{E}\left[H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A,\mathbf{X},R=k\right]=0$
@@
$=\mathbb{E}\bigg{[}S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k\mid\mathbf{X})}\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}$
×$\qquad\times\left\{H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\bigg{]}.$
@@
$\frac{\partial}{\partial\epsilon}\theta^{0}(t,a)\bigg{|}_{\epsilon=0}$
$=\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\mathbb{E}\bigg{\{}S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k\mid\mathbf{X})}$
×$\quad\quad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(y\mid\mathbf{X})G^{k}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})~{}\bigg{|}~{}\mathbf{X}\bigg{\}}\bigg{]}$
$=\mathbb{E}\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\{S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)\}\dot{\ell}(\mathcal{O})\right]-\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=0)}\frac{\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=k\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
×$\quad\quad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(y\mid\mathbf{X})G^{k}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\bigg{]}.$
@@
$\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})$
$=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=0)\mathbb{P}(R=k\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
×$\qquad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right].$
@@ D.2.2 Regularity conditions and RAL of the source-site estimator
If Conditions D.6D.7 hold, with $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$, the CCOD holds, and Condition D.8 also holds, then $\widehat{\theta}^{k,0}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*k,0}_{t,a})+o_{p}(n^{-1/2})$. In particular, $n^{1/2}(\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a))$ then convergences in distribution to a normal random variable with mean zero and variance $\sigma^{2}=\mathbb{P}[(\varphi^{*k,0}_{t,a})^{2}]$. If Conditions D.9 and D.10 also hold, then
@@ D.2.3 Lemmata for the source-site estimator
$\widehat{\theta}_{n}^{k}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*k,0}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right].$ (7)
@@
Under Condition D.7, there exists a universal constant $C=C(\eta)$ such that for each $k$, $m$, $n$, $t$, and $a$,
@@
$B^{k}(\mathcal{V}_{m})$ $=\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
×$\quad\times\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right],$
@@
The expression of $\widehat{B}^{k}_{m}(\mathcal{V}_{m})-B^{k}_{\infty}(\mathcal{V}_{m})$ is exactly the same as the Lemma 3 in Westling et al. [48], while we only need to replace the corresponding nuisance functions by the site-$k$ version here, so the detail is omitted. By the triangle inequality, we have $\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}-\varphi^{k,0}_{\infty,t,a}]^{2}\leq\left\{\sum_{j=1}^{4}\{\mathbb{P}[(U^{k}_{j,n,m,t,a})^{2}]\}^{1/2}\right\}^{2}$. Therefore, under Assumption 2.3 and Condition D.7, there exists a universal constant $C=C(\eta)$ such that the result in the statement holds. Thus, the proof is completed. ∎
@@
If Conditions D.6D.7 hold, $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}^{1/2}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a_{0}}\right]=o_{p}(n^{-1/2})$. If Conditions D.9 holds as well,
@@
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}S^{k}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\omega^{k,0}_{\infty}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})\pi^{k}_{\infty}(a\mid\mathbf{X})}-1\right\}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$
@@
$\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\{S^{0}_{\infty}(t\mid a,\mathbf{X})-S^{0}(t\mid a,\mathbf{X})\}+\frac{q^{k}(\mathbf{X})}{\mathbb{P}(R=k)}\omega^{k,0}_{\infty}(\mathbf{X})S^{k}_{\infty}(t\mid a,\mathbf{X})\frac{\pi^{k}(a\mid\mathbf{X})}{\pi^{k}_{\infty}(a\mid\mathbf{X})}$
×$\qquad\qquad\times\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})G^{k}(y\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}$
$=\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\{S^{0}_{\infty}(t\mid a,\mathbf{X})-S^{0}(t\mid a,\mathbf{X})\}+\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\omega^{k,0}_{\infty}(\mathbf{X})}{\omega^{k,0}(\mathbf{X})}S^{k}_{\infty}(t\mid a,\mathbf{X})\frac{\pi^{k}(a\mid\mathbf{X})}{\pi^{k}_{\infty}(a\mid\mathbf{X})}$
×$\qquad\qquad\times\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})G^{k}(y\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$